### v1

# Culturally transmitted color categories in LLMs reflect a learning bias toward efficient compression

Abstract
Converging evidence suggests that systems of semantic categories across human languages achieve near-optimal compression via the Information Bottleneck (IB) complexity-accuracy principle. Large language models (LLMs) are not trained for this objective, which raises the question: are LLMs capable of evolving efficient human-like semantic systems? To address this question, we focus on the domain of color as a key testbed of cognitive theories of categorization and replicate with LLMs (Gemini 2.0-flash and Llama 3.3-70B-Instruct) two influential human behavioral studies. First, we conduct an English color-naming study, showing that Gemini aligns well with the naming patterns of native English speakers and achieves a significantly high IB-efficiency score, while Llama exhibits an efficient but lower complexity system compared to English. Second, to test whether LLMs simply mimic patterns in their training data or actually exhibit a human-like inductive bias toward IB-efficiency, we simulate cultural evolution of pseudo color-naming systems in LLMs via iterated in-context language learning. We find that akin to humans, LLMs iteratively restructure initially random systems towards greater IB-efficiency and increased alignment with patterns observed across the world’s languages. These findings demonstrate that LLMs are capable of evolving perceptually grounded, human-like semantic systems, driven by the same fundamental principle that governs semantic efficiency across human languages.
1 Introducton
As large language models (LLMs) are increasingly deployed in real-world everyday settings, it is crucial to understand how their learning biases and representational capacities align with our own. We investigate this by focusing on a key aspect of human intelligence: the ability to organize information into semantic categories Croft (2002); Rosch (2002); Koch (2008); Boster (2005); Majid (2015); Malt and Majid (2013). This phenomenon reveals two major challenges for AI. First, semantic categories exhibit both universal patterns and cross-language differences Berlin and Kay (1969); Berlin (1992); Croft (2002), which LLMs must navigate. Second, LLMs are not grounded in the rich physical and social environment that humans are, and it is unclear how these differences affect their ability to learn human-aligned semantic categories Rosch (1975); Labov (1973). Therefore, in order to understand whether LLMs can efficiently communicate with people and adapt to changing environments and communicative needs, it is crucial to study whether LLMs are capable of structuring meaning according to the same principles that guide humans. Here, we address this challenge and propose a novel cognitively-motivated framework for studying semantic systems in LLMs. We build on the theoretical framework of Zaslavsky et al. (2018), which argues that languages efficiently compress meanings into words by optimizing the Information Bottleneck (IB) principle Tishby et al. (1999), instantiated as a tradeoff between the informational complexity and communicative accuracy of the lexicon. This framework has broad empirical support across languages Zaslavsky et al. (2018, 2019, 2021); Mollica et al. (2021); Zaslavsky et al. (2022). Furthermore,  Imel et al. (2025) recently showed that a drive for IB-efficiency may be present in the individual inductive biases of human learners. LLMs, however, are not trained with respect to the IB objective, which raises the question: are LLMs capable of evolving efficient human-like semantic systems?
To address this question, we focus on color, a key test case for categorization theories in cognitive science with rare cross-cultural human data Berlin and Kay (1969); Regier et al. (2007, 2015); Gibson et al. (2019), and replicate with LLMs two influential human behavioral experiments: an English color naming experiment Lindsey and Brown (2014), designed to assess the efficiency and human-alignment of the color naming systems of LLMs; and an iterated language learning (ILL) experiment Xu et al. (2013), designed to probe the implicit inductive biases of LLMs by simulating cultural transmission Kirby et al. (2014). For the latter, we develop iterated in-context language learning (IICLL). Because IICLL requires strong in-context capabilities that are too demanding for most open-weights models, we focus on two prominent LLMs that are capable of performing the task: Gemini 2.0-flash Google (2025), a top proprietary multi-modal model, and Llama 3.3-70B-Instruct AI@Meta (2024), a leading open-weights model.
We show that (1) Gemini aligns well with the English color naming system and achieves high IB-efficiency, while Llama, the smaller model, exhibits a highly efficient but lower complexity system; and (2) similar to humans, both LLMs iteratively restructure initially random systems towards greater IB-efficiency and increased human-alignment, suggesting that they do not simply imitate patterns in their training data but are intrinsically guided by a drive toward IB-efficiency. These findings demonstrate that LLMs are capable of evolving perceptually grounded, human-like semantic systems, driven at least in part by the same efficiency principle that underlies human languages.
2 Background
The Information Bottleneck (IB) color naming model.  We apply the theoretical framework from Zaslavsky et al. (2018), which models communication as a speaker mapping a mental representation ($m$) of a world state ($u$) to a word ($w$) via an encoder $q(w|m)$. In the case of color, the world states $\mathcal{U}$ are given by a set of target colors (Figure˜1C). To account for perceptual noise, each $m$ is modeled as a Gaussian distribution over the perceptual CIELAB color space centered around a target color Zaslavsky et al. (2018). To communicate efficiently, the speaker and listener optimize the Information Bottleneck (IB) tradeoff Tishby et al. (1999) between minimizing the complexity of their lexicon, $I_{q}(M;W)$, and maximizing its accuracy, $I_{q}(W;U)$. An optimal semantic system minimizes the IB objective:
$\mathcal{F}_{\beta}[q]=I_{q}(M;W)-\beta I_{q}(W;U),$ (1)
where $\beta\geq 1$ controls the tradeoff, and the solutions to this optimization problem define the IB bound of efficiency. Figure˜1F shows the IB bound for color naming from Zaslavsky et al. (2018) together with a reproduction of their results showing that color naming systems across languages achieve near-optimal tradeoffs. The empirical basis for this analysis is the World Color Survey (WCS) dataset Cook et al. (2005), containing naming data from 110 languages of non-industrialized societies elicited via the task illustrated in Figure˜1A, with respect to the WCS grid shown in Figure˜1C. Our analysis also includes English color naming data from Lindsey and Brown (2014) (Figure˜1D).
Iterated language learning (ILL) of color naming systems.  In iterated language learning (ILL), human participants form “generations” in which they learn a language from a sample of the previous generation’s output and then produce a new language system $(L_{t})$ for the next. This process has been used to reveal how learners’ inductive biases emerge from generalizing on limited data Griffiths and Kalish (2007); Griffiths et al. (2008). We adapt this paradigm for LLMs, yielding iterated in-context language learning (IICLL) (see Figure˜1B and Section˜3.2). The empirical comparison for our IICLL study is the human ILL data from Xu et al. (2013), in which participants learned and transmitted novel color naming systems across chains of generations. Each chain was initialized with a random system, and at each generation, participants were trained on a small set of color-pseudoword pairs before labeling the entire grid. Over time, chains became increasingly regular and more similar to systems in the WCS dataset Xu et al. (2013). More recently, Imel et al. (2025) showed that these chains also converge to highly efficient systems along the IB bound (see Figure˜1F).
Figure 1: A. Color naming task with humans (top) and LLMs (bottom). Multi-modal LLMs can observe colors either via text or images. B. The IICLL paradigm. At each generation $t$, an LLM is prompted with a small dataset, $d_{t-1}$, consisting of colors with pseudo labels sampled from the previous generation’s language, $L_{t-1}$. With $d_{t-1}$ in context, the LLM performs the naming task for the full meaning space. C. The WCS color naming grid. D-E. Color naming systems of English speakers and LLMs. Each system is plotted against the WCS grid; each chip is colored by the color-centroid of its modal category. F-G. Humans and LLM on the information plane from Zaslavsky et al. (2018), where the black curve is the IB theoretical bound. F. Gemini’s English color naming system achieves a similar tradeoff as that of native English speakers, while Llama exhibits a lower complexity system. G. IICLL with LLMs converges to near-optimal solutions (colored dots) within the same range of human IL chains and WCS languages (shown in F). Small black dots correspond to random initialization of chains with varying number of categories, $k\in\{2,3,4,5,6,14\}$. Thin lines correspond to IL trajectories. H. Over generations of IICLL, LLM systems become more efficient (left), more aligned to IB optima (middle), and more aligned to WCS languages (right). Colored curves represent averages across initializations and conditions, with the shaded region indicating the 95% CI.
3 Experiments
We aim to test whether LLMs are capable of acquiring a human-like inductive bias toward IB efficiency, even though they are not explicitly trained for this objective function. To this end, we implement with LLMs a color naming experiment and an IICLL experiment, designed to assess the color category systems and inductive learning biases implicit in LLMs. We then compare the LLM data with the corresponding human data through the lens of the IB framework. Specifically, we focus on three evaluation measures: (1) efficiency loss, measuring the LLM’s deviation from IB-optimality, i.e., with respect to Equation˜1; (2) IB-alignment, measuring the structural similarity between an LLM system and its nearest IB-optimal system; and (3) human-alignment, measuring the structural similarity between an LLM system and human systems (from the WCS or English). Because our IICLL task requires strong in-context capabilities that are too demanding for most open-weights models we tested, we focus on two prominent LLMs that are capable of performing the task: Gemini 2.0-flash Google (2025), a leading industry-level multi-modal model, and Llama 3.3-70B-Instruct AI@Meta (2024), a leading open-weights text-based model.
3.1 English Color Naming with LLMs
We begin with a color naming study (Figure˜1A) to assess the semantic alignment and communicative efficiency of LLMs with respect to native English speakers. Each LLM was prompted to label all $330$ colors shown in Figure˜1C, one at a time in a randomized order. These stimuli were presented either as sRGB coordinates (txt input), which both models can process, or as colored image patches (img input), which only Gemini can process. In both cases, the models were instructed to use the set of English terms from the free-naming experiment of Lindsey and Brown (2014).
The resulting LLM color naming systems are shown in Figure˜1E, their complexity-accuracy tradeoffs are shown in Figure˜1F relative to the IB bound, and their efficiency and alignment scores are shown in Table˜1. Gemini, using either txt or img inputs, demonstrates the highest alignment with the human English system and reaches human-like IB tradeoffs (Figure˜1F). Llama produces a highly efficient system but with lower complexity compared to English. This suggests that both LLMs employ highly efficient color naming systems, but the model’s size may impact its complexity range and in turn, its ability to align with humans. Finally, to ensure these results robustly reflect the LLM behavior, rather than brittle memorization, we also replicated the same results with added input noise which is imperceptible.
Table 1: Evaluation of the efficiency loss (lower values are better) and alignment (higher values are better) of LLMs. Results show averaged over ten random seeds ($\pm$ STD).
Model Efficiency Loss IB-alignment English-alignment
Llama-txt 0.68 (0.03) 0.45(0.02)0.45~(0.02) 0.55(0.02)0.55~(0.02)
Gemini-img 0.89(0.03)0.89~(0.03) 0.52(0.01)0.52~(0.01) 0.69 (0.01)
Gemini-txt 0.79 (0.03) 0.53 (0.01) 0.65 (0.01)
3.2 Iterated in-context language learning (IICLL)
To further understand the inductive learning biases of LLMs, beyond what they may have seen during training, we study how LLMs can evolve category systems on their own through a process of cultural evolution. We do so by replicating as closely as possible the ILL color naming experiment of Xu et al. (2013), while adapting it to LLMs. To this end, we develop a method which we call iterated in-context language learning (IICLL, Figure˜1B), consisting of a chain of generations simulated by an LLM. In each generation, the LLM learns in-context an artificial color naming system produced by the previous generation (or randomly initialized for the first generation) from a small random sample of color-signal pairs. Importantly, the signals are not real words but rather pseudowords, designed to avoid a bias toward existing languages. Given this in-context learning phase, the model is then prompted to perform the full naming task (Figure˜1A), generalizing the meanings of the pseudowords to the full color space of Figure˜1C. This reveals the model’s implicit inductive biases, which are then amplified with generations.
Figure˜1G shows the IICLL trajectories of both models. Over generations, the LLMs approach the IB bound, becoming more efficient (Figure˜1H, left) and more aligned with the IB-optimal systems (Figure˜1H, middle), above and beyond a strong counterfactual baseline of hypothetical variants. Interestingly, Llama appears to be confined to a lower complexity region compared to human languages, while Gemini is capable of spanning the full complexity range observed in humans (comparing the WCS and IL data in Figure˜1F with the LLM end-points in Figure˜1G). Finally, Figure˜1H shows that Gemini-txt is more efficient and more human-aligned compared to Gemini-img, although Gemini-img seems to track more closely the trends observed in the human iterated learning (IL) data. This suggests that the input representation may have an important role in how LLMs generalize semantic knowledge and categories. One possibility, for example, is that it might be easier for Gemini to process textual RGB inputs, leading it to more efficient solutions, while visual inputs leads it to more human-like behavior.
4 Conclusions
This work combines a theory-driven approach, based on the IB principle, with cognitively-motivated experimental methods, based on color naming and iterated language learning, to study whether LLMs can acquire a human-like inductive bias toward optimally-compressed semantic representations without being trained for this objective. We found that LLMs can align well with human color naming systems and tend to restructure randomly-initialized artificial category systems toward greater IB-efficiency and human-alignment with human naming systems. This demonstrates that LLMs are capable of evolving perceptually grounded, human-like semantic systems, guided by the same IB-efficiency principle that underlies human languages. Our findings also suggest that model size and input representations may be key factors in determining the level of complexity and human-alignment of LLMs. Finally, our work lays out the foundations for further evaluations of this capacity across more semantic domains and models.
5 Acknowledgments
This work was support by a Google Cloud Platform Credit Award.